OpenStudy (anonymous):

I'm trying to use partial fractions to find the integral of 1/(3x-3x^2) dx ...

OpenStudy (anonymous):

So what I have so far is: I factored the denominator into (x)(-3x+3) So A/(x) + B/(-3x+3) = 1 so A(-3x +3) + B(x) = 1 But now I don't know what to do ...

OpenStudy (anonymous):

can we write \[\int\limits_{?}^{?}\frac{ dx }{ 3x - x^2 } = \int\limits_{?}^{?}\frac{ dx }{ 3x(1-x) }\] ?

OpenStudy (anonymous):

Sure, so then it would be A/3x + B/(1-x) so A(1-x) + B(3x) = 1 But same problem ... what now?

OpenStudy (anonymous):

i will provide you easy solution \[\frac{ 1 }{ 3 }\int\limits_{}^{}\frac{ 1 }{ x(1-x) }*dx\] add and subtract x in the numerator \[\frac{ 1 }{ 3 } \int\limits_{}^{} \frac{ x+1-x }{x(1-x) }*dx\] do u follow these steps?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now can u separate the terms?

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

can we write above as \[\frac{ 1 }{ 3 }\int\limits_{}^{} \frac{ x }{ x(1-x) } + \frac{ 1-x }{ x(1-x) }\] ?

OpenStudy (anonymous):

OH. I see.

OpenStudy (anonymous):

can u do the rest part?

OpenStudy (anonymous):

i'm not really sure where you're going with this is the thing.

OpenStudy (anonymous):

after cancelling the like terms, we will get \[\frac{ 1 }{ 3 }\int\limits_{}^{} (\frac{ 1 }{ 1-x }+ \frac{ 1 }{ x })\] i think the integration has left only hope u can do this now

OpenStudy (anonymous):

Yes! OK. I did my process differently but got this in the end as well.

OpenStudy (anonymous):

if u want to solve the ques by your way then i can tell you that way too i just told u different method

OpenStudy (anonymous):

No, no, I get it. I like your method, also.

OpenStudy (anonymous):

One more question: I'm getting 1/3 ln (p) + ln (1-p) But I'm supposed to be getting 1/3 ln (p/1-p) Where did I go wrong?

OpenStudy (anonymous):

are u getting A= 1/3 and B = 1/3 ?

OpenStudy (anonymous):

No, I'm getting A = 1/3 and B = 1

OpenStudy (anonymous):

because -3A + B = 0 and A = 1/3 so B should be 1 right?

OpenStudy (anonymous):

sorry my fault B should be 1

OpenStudy (anonymous):

K

OpenStudy (anonymous):

so u will be getting \[\int\limits_{}^{}\frac{ A }{ x } + \frac{ B }{ (-3x+3) }\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Use A=1/3 and B=1 and also make sure to take 3 out of the denominator

OpenStudy (anonymous):

Right, I did use those values, which gives me 1/3 * (1/p + 1/(1-p))

OpenStudy (anonymous):

and the integral of that is 1/3[ ln(p) + ln(1-p) ], right? But that's not the answer I'm supposed to be getting.

OpenStudy (anonymous):

there would be "-ve" sign u forgot to take the derivate of the terms inside natural log {remember chain rule }

OpenStudy (anonymous):

Oh. So 1/3 [ ln(p) - ln(1-p)?

OpenStudy (anonymous):

yah now use the proprty of log

OpenStudy (anonymous):

I'm supposed to get 1/3 [ ln (p/1-p) ], but I don't know why you divide by 1-p, that doesn't make sense.

OpenStudy (anonymous):

OH! BECAUSE OF THE PROPERTIEES OF LOGS!

OpenStudy (anonymous):

yeah property of natural log is 1) lnA - lnB = ln(A/B) 2) lnA +lnB = ln(A*B)

OpenStudy (anonymous):

YAY!!! OK thank you so much I understand this now :D

OpenStudy (anonymous):

:-P